# Operatorial formulation of crimo-taxis phenomena in a street

### v1

@@ 2 Operatorial framework
$\dot{a}_{1,\alpha}=\textrm{i}\left(-\widetilde{\omega}_{1,\alpha}a_{1,\alpha}+2\nu_{3}a_{3,\alpha}+\widetilde{\lambda}_{1,\alpha}a_{2,\alpha}+\widetilde{\mu}_{1}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}\right.$ (12)
$\qquad\left.+\widetilde{\nu_{1}}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1}}{\Delta x^{2}}+\widetilde{\nu_{3}}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right),$
$\dot{a}_{2,\alpha}=\textrm{i}\left(-\widetilde{\omega}_{2,\alpha}a_{2,\alpha}+\widetilde{\lambda}_{1,\alpha}a_{1,\alpha}+\widetilde{\lambda}_{2,\alpha}a_{3,\alpha}+\widetilde{\mu}_{2}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1}}{\Delta x^{2}}\right.$
$\qquad\left.+\widetilde{\nu_{1}}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}+\widetilde{\nu_{2}}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right),$
$\dot{a}_{3,\alpha}=\textrm{i}\left(-\widetilde{\omega}_{3,\alpha}a_{3,\alpha}+2\nu_{3}a_{1,\alpha}+\widetilde{\lambda}_{2,\alpha}a_{2,\alpha}+\widetilde{\mu}_{3}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right.$
$\qquad\left.+\widetilde{\nu_{2}}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1})}{\Delta x^{2}}+\widetilde{\nu_{3}}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}\right),$
@@
Particular emphasis is placed on the last three terms of each equation in the system (12), which can be interpreted as the spatial discretization of the second-order spatial derivative (with step size $\Delta x$). These terms represent the self-diffusion and cross-diffusion, respectively. Therefore, an operatorial model of a lattice can be thought of as the spatial discretization of a system of partial differential equations of reaction-diffusion type.
@@ 2.1 Analytical solution
The unknowns in (11) are operators represented by $2^{3N}\times 2^{3N}$ matrices with complex entries; whereupon, at least in principle, we have to solve a system of $3N\cdot 2^{6N}$ scalar ordinary differential equations in the complex domain. Nevertheless, because of the linearity of equation (11), we are able to reduce drastically the computational complexity. In fact, introducing the formal column vector
@@
(where the symbol T stands for the transposition operator) and a suitable $3N\times 3N$ real matrix $\Gamma$ (whose entries, once we fix $N$, can be constructed from (11)), the evolution equations for the annihilation operators can be written in the compact form as
$\frac{d\mathbf{A}}{dt}=\textrm{i}\Gamma\,\mathbf{A}.$ (14)
Thence, the formal solution of the system (14) is expressed by
@@
Here, we denote by $n_{i,\alpha}^{0}$ the density of population $\mathcal{P}_{i}$ in the cell $\alpha$ at the initial time $t_{0}=0$, and by $\mathbf{n}^{0}$ the corresponding initial values, say
$\mathbf{n}^{0}=\left({n_{1}^{0}},\ldots,{n_{3N}^{0}}\right),$
where $n_{k}^{0}$ is the initial density of $j$th population at the cell $\alpha$, with $k=(j-1)N+\alpha$.
Denoting with $B_{j,k}$ the generic entry of the $3N\times 3N$ matrix $B(t)$, and
$n_{j,\alpha}(t)=\langle\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}},a^{\dagger}_{j,\alpha}(t)a_{j,\alpha}(t)\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}\rangle,\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N,$
it is easy to derive, by using the CAR (3), the formula giving the mean values of the number operators at time $t$ [24]
$n_{j,\alpha}(t)=\sum_{\beta=1}^{3N}\left|B_{(j-1)N+\alpha,\beta}(t)\right|^{2}n_{\beta}^{0},\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N.$ (16)
The quadratic form of the Hamiltonian has an immediate consequence: the dynamic behavior that we can obtain in each cell is at most quasiperiodic. Furthermore, due to the relation
@@
the Hamiltonian possesses a first integral, expressing the fact that the sum of the densities of the three populations all over the cells of the domain is constant in time.
More complex outcomes (not necessarily quasiperiodic) could be recovered including in the Hamiltonian terms more than quadratic [9, 31]. Unfortunately, in such case we would be forced to solve numerically a very huge number of differential equations. Another strategy, with the goal of enriching the dynamics without rendering the problem computationally hard, has been recently introduced in [7, 23, 28, 32, 33], where interesting dynamics evolution is obtained still retaining a quadratic and time-independent Hamiltonian.
3 ($\mathcal{H},\rho$)–induced dynamics
The aim of the $(\mathcal{H},\rho)$–induced dynamics approach is to modify the standard Heisenberg dynamics by introducing some rules able to account for some effects that cannot be embedded in the definition of the Hamiltonian, unless we do not assume explicitly a time dependent Hamiltonian or consider an open quantum system including a reservoir, with consequent increase of technical difficulties (see [7]).
The concept of rule is nothing more than a law that modifies periodically some of the values of the parameters involved in the Hamiltonian as a consequence of the evolution of the system, so that the model is somehow able to adjust itself during the time evolution; essentially, this mathematical trick allows the actors to modify their attitudes and the strength of interactions according to the evolution of their state.
Hereafter, we sketch briefly the steps of the procedure (see [7] and references therein for further details):
  1. 1.
    divide the time interval $[0,T]$ in $n$ subintervals of length $\tau$;
  2. 2.
    let $k=1$, and consider a Hamiltonian operator $\mathcal{H}^{(k)}$;
  3. 3.
    using Heisenberg view, compute, in the time interval $[(k-1)\tau,k\tau]$, the evolution of ladder operators, whereupon, choosing an initial condition for the mean values of the number operators, obtain the time evolution of the latter (our observables);
  4. 4.
    according to the values of the observables at time $k\tau$, or to their variations in the time interval $[(k-1)\tau,k\tau]$, modify some of the parameters involved in $\mathcal{H}^{(k)}$;
  5. 5.
    a new Hamiltonian operator $\mathcal{H}^{(k+1)}$, having the same functional form as $\mathcal{H}^{(k)}$, but (in general) with different values of (some of) the involved parameters, is obtained;
  6. 6.
    increment by 1 the value of $k$;
  7. 7.
    if $k<n$ go to step 3.
From a mathematical point of view, the rule is just a map from $\mathbb{R}^{p}$ into $\mathbb{R}^{p}$ acting on the space of the $p$ parameters involved in the Hamiltonian. The global evolution is governed by a sequence of similar Hamiltonian operators, and the parameters entering the model can be considered stepwise (in time) constant. The complete evolution in the interval $[0,T]$ is obtained by glueing the local evolutions in each subinterval.
By using the $(\mathcal{H},\rho)$–induced dynamics approach, it is implicitly assuming to have a time dependent Hamiltonian. However, the time dependence is, in our case, of a very special form: in each interval [$(k-1)\tau,k\tau[$ the Hamiltonian does not depend on time, but in $k\tau$ some changes may occur, according to how the system is evolving. A comparison of this approach with that related to an explicitly time dependent Hamiltonian is discussed in [7].
In the model investigated in this paper, after fixing the value of $\tau$ (i.e., the length of the subintervals), we update at instants $k\tau$ ($k$ integer) the parameters as specified below.
Let us define
@@
where $\delta_{j,\alpha}^{(k)}$ denotes the variation of the mean values of the number operators associated to each subgroup, i.e., how the subgroup densities vary at each step, and $\delta_{\alpha}^{(k)}$ is the difference between the products of the mean values of the number operators, i.e., it represents the variation of the interaction between ordinary citizens and drug addicts.
Now, let us consider the following rules for the three families of inertia and diffusion parameters, respectively, that come into play in the Hamiltonian by updating them at fixed times $k\tau$ ($k=1,2,\ldots)$, say
@@
$\mu_{3,\alpha}\,\rightarrow\,\mu_{3,\alpha}\left(1-\delta_{\alpha}^{(k)}\right).$ (19c)
@@
Some comments about the rule are in order. The relation (18a) indicates that the inertia parameter of the population $\mathcal{P}_{1}$ in the cell $\alpha$ increases (or decreases) when the difference between the local densities of police and drug addicts in the subinterval of length $\tau$ increases (or decreases). This suggests that a population increasing its inertia parameter tends to lower its tendency to change. The rationale of this rule is that an increase of the local density of this difference in the cell $\alpha$ has the effect of inducing a lower tendency to change (the police local density is greater than the one of drug addicts, so the population would feel safe), whereas a decrease induces the population to be less conservative (the local density of drug addicts is greater than the police one, so the population would feel less safe).
@@
The relation (18b) states that the inertia parameter of the population $\mathcal{P}_{2}$ in the cell $\alpha$ increases (or decreases) as the difference between the local densities of the ordinary citizens and police increases (or decreases). This rule takes into account that an increase of the local density of this difference in the cell $\alpha$ has the effect of inducing a lower tendency to change (the population local density is greater than police one, therefore drug addicts continue to commit crime), whereas a decrease induces the drug addicts to be less conservative (the police local density is greater than the one of ordinary citizens, so that the drug addicts will tend to move to other areas to commit crimes).
@@
The rule for inertia parameters of law enforcement personnel, as expressed in (18c), incorporates the concept of social alarm introduced by Epstein [1]. The inertia parameters are influenced by the sum of local densities of interactions between the ordinary citizens and drug addicts, alongside the local density of police. This social alarm is crucial for illustrating scenarios where law enforcement intervention is necessary to restore order.
@@
The relation (19a) indicates that the diffusion parameter of the subgroup $\mathcal{P}_{1}$ in the cell $\alpha$ increases (or decreases) when the difference between the drug addicts and police local densities increases (or decreases). The idea behind it is that citizens must diffuse to neighboring territories when there is a high crime rate (the difference is positive), but if the police are sufficient in number to maintain order (the difference is negative), citizens can remain in their safe place.
@@
The relation (19b) shows that the diffusion parameter of the drug addicts $\mathcal{P}_{2}$ in the cell $\alpha$ increases (or decreases) as the difference between the police and citizen local densities increases (or decreases). Drug addicts are less likely to dissolve when citizens density is high and police one is low, as this encourages continued criminal activity. In fact, with a high police presence, drug addicts may relocate to seek opportunities for crime elsewhere.
@@
Finally, the relation (19c) reveals that the diffusion parameter of police $\mathcal{P}_{3}$ in the cell $\alpha$ increases (or decreases) as interactions between the drug addicts and citizens decrease (or increase). This implies that police should spread to other cells when social alarm in the cell $\alpha$ is low.
@@ 4 Numerical results
In this Section, we present some numerical simulations for the model (11).
In the first part, we neglect spatial inhomogeneities, i.e., all the interactions occur just in a single cell, whereas, the second one concerns with the model investigated in a one–dimensional spatial domain made of $N=50$ cells. In both situations, we compare the numerical results obtained by the standard Heisenberg dynamics and those coming from the superposition of the Heisenberg dynamics with the rules above described.
@@ 4.1 Model in one cell
In the standard Heisenberg dynamics, we see in Figure 2 that the evolution exhibits a never ending oscillatory trend, as expected. On the contrary, when the rule (18) is used, the amplitude of the oscillations tends to decrease and the mean densities of the three subgroups tends to reach an asymptotic equilibrium (see Figure 2). Furthermore, from Figure 2, an interpretation can be that law enforcement personnel are able to keep better drug users/dealers under control. Note also that the time needed to get a stationary behavior after the transient behavior depends on the values of the parameters.
Figure 2: Time evolution of the mean values of the number operators of the three subgroups. Only one cell is considered. No rule (Figure 2), and rule (18) with $\tau=4$ (Figure 2). The initial conditions in both cases are: $n_{1}^{0}=0.8$, $n_{2}^{0}=0.08$, $n_{3}^{0}=0.04$. The inertia and interaction parameters are $\omega_{1}=0.5$, $\omega_{2}=0.3$, $\omega_{3}=0.2$, and $\lambda_{1}=0.04$, $\lambda_{2}=0.025$, respectively.

### v2

@@ 2 Operatorial framework
$\dot{a}_{1,\alpha}(t)$ $=\textrm{i}\left(-\widetilde{\omega}_{1,\alpha}a_{1,\alpha}+2\nu_{3}a_{3,\alpha}+\widetilde{\lambda}_{1,\alpha}a_{2,\alpha}+\widetilde{\mu}_{1}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}\right.$ (12)
$+\left.\widetilde{\nu_{1}}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1}}{\Delta x^{2}}+\widetilde{\nu_{3}}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right),$
$\dot{a}_{2,\alpha}(t)$ $=\textrm{i}\left(-\widetilde{\omega}_{2,\alpha}a_{2,\alpha}+\widetilde{\lambda}_{1,\alpha}a_{1,\alpha}+\widetilde{\lambda}_{2,\alpha}a_{3,\alpha}+\widetilde{\mu}_{2}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1}}{\Delta x^{2}}\right.$
$+\left.\widetilde{\nu_{1}}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}+\widetilde{\nu_{2}}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right),$
$\dot{a}_{3,\alpha}(t)$ $=\textrm{i}\left(-\widetilde{\omega}_{3,\alpha}a_{3,\alpha}+2\nu_{3}a_{1,\alpha}+\widetilde{\lambda}_{2,\alpha}a_{2,\alpha}+\widetilde{\mu}_{3}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right.$
$+\left.\widetilde{\nu_{2}}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1})}{\Delta x^{2}}+\widetilde{\nu_{3}}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}\right),$
@@
Particular emphasis is placed on the last three terms of each equation in the system (12), which can be interpreted as the spatial discretizations of the second-order spatial derivative (with step size $\Delta x$). These terms represent the self-diffusion and cross-diffusion, respectively. Therefore, an operatorial model of a lattice can be thought of as the spatial discretization of a system of partial differential equations of reaction-diffusion type.
@@ 2.1 Analytical solution
In this Subsection, we provide an explicit form for the unknowns involved in the system (11); these are operators expressed by $2^{3N}\times 2^{3N}$ matrices, thus implying that, without any further simplification, $3N\cdot 2^{6N}$ scalar ordinary differential equations have to be solved in a complex domain. However, since the system (11) is made of linear differential equations, the computational complexity can be drastically reduced. Let us collect the $3N$ operators $a_{j,\alpha}$ in the column vector $\mathbf{A}$, say
@@
(where T denotes transposition), and consider a $3N\times 3N$ matrix $\Gamma$ (whose real entries, after the number of agents $N$ is fixed, can be derived from equations (11)); then, the ordinary differential equations ruling the evolution of the annihilation operators can be read in the following compact form
$\frac{d\mathbf{A}}{dt}=\textrm{i}\Gamma\,\mathbf{A},$ (14)
whose analytical solution is expressed by
@@
Here, we denote with $n_{j,\alpha}^{0}$ the $j$th subgroup density in the cell $\alpha$ at the initial time $t_{0}=0$, and with $\mathbf{n}^{0}$ the corresponding initial value, say
$\mathbf{n}^{0}=\left({n_{1}^{0}},\ldots,{n_{3N}^{0}}\right)^{T},$
where $n_{k}^{0}$ is the initial density of $j$th subgroup at the cell $\alpha$, with $\allowbreak k=(j-1)N+\alpha$.
Given the generic entry $B_{j,k}$ of the $3N\times 3N$ matrix $B(t)$, using the relation
$n_{j,\alpha}(t)=\langle\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}},a^{\dagger}_{j,\alpha}(t)a_{j,\alpha}(t)\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}\rangle,\qquad j=1,2,3,\qquad\alpha=1,\ldots,N,$
and the CAR (3), the mean values of the number operators at time $t$ can be easily recovered [23]:
$n_{j,\alpha}(t)=\sum_{\beta=1}^{3N}\left|B_{(j-1)N+\alpha,\beta}(t)\right|^{2}n_{\beta}^{0},\qquad j=1,2,3,\qquad\alpha=1,\ldots,N.$ (16)
The approach above described reveals really useful but has some limitations, as detailed in the following Remark.
Remark 2.
  • A first integral is possessed by the Hamiltonian $\mathcal{H}$, say
@@
    i.e., sum of the three subgroup densities all over the cells of the entire domain remains constant in time.
  • The Hamiltonian has a quadratic structure: as a consequence, the dynamics in each cell is at most quasiperiodic.
An alternative strategy to obtain different (especially not quasiperiodic) dynamical behaviors could be that of incorporating higher order terms in the Hamiltonian, ora considering an open quantum system [8, 31]. Unfortunately, the computational difficulties will increase because a very huge number of differential equations must be numerically solved (see [27]). A recent approach, with the goal of enriching the dynamics without increasing the computational complexity, has been introduced in [27, 22, 28, 32, 33], where non-oscillatory trends can be recovered still keeping an Hamiltonian with a time-independent and quadratic structure.
3 The ($\mathcal{H},\rho$)–induced dynamics approach
The aim of the $(\mathcal{H},\rho)$–induced dynamics approach is to avoid the quasiperiodic behavior, that is due to the classical Heisenberg view, by including in the Hamiltonian some effects: this is done using suitable rules. More in detail, a rule is a law that changes, at fixed instants, some of the values of the parameters entering the Hamiltonian, so that the model is somehow able to adapt itself according to the time evolution of the system. From a mathematical viewpoint, a rule is a mapping $\rho\colon\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ working on the $p$ parameters included in the Hamiltonian. Essentially, the main advantage of this mathematical expedient is to allow the agents to modify, according to their evolution, their attitudes and interaction rates.
Within the framework of $(\mathcal{H},\rho)$–induced dynamics (see [27] for a complete description), the values of some parameters involved in the Hamiltonian are stepwise constant in time, as periodically they are adjusted according to the evolution of the system itself. From a computational point of view, this implies that the complete evolution of the system in the time interval $[0,T]$ is obtained by glueing the local evolutions in a finite number of subintervals. In each subinterval the Hamiltonian is time-independent, but globally we have a time-dependent Hamiltonian (see [27] for a comparison of this approach with the general case of a time-dependent Hamiltonian).
According to this method, we fix the length $\tau$ of the subintervals, and, then, update at each instant $k\tau$ ($k$ integer) the values of the parameters as specified below. At first, let us introduce the quantities
@@ 3 The ($\mathcal{H},\rho$)–induced dynamics approach
where $\delta_{j,\alpha}^{(k)}$ denotes the variation of the mean values of the number operators in the cell $\alpha$ associated to each subgroup $\mathcal{P}_{j}$, i.e., how the subgroup local densities vary at each step, and $\delta_{\alpha}^{(k)}$ is the difference between the products of the mean values of the number operators associated to the subgroups $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, i.e., it represents the variation at each step of the interactions between ordinary citizens and drug addicts.
More in detail, let us consider the following rules for the three families of inertia and diffusion parameters, respectively, say
@@
$\mu_{3,\alpha}\,\rightarrow\,\mu_{3,\alpha}\left(1-\delta_{\alpha}^{(k)}\right),$ (19c)
holding for all $\alpha=1,\ldots,N$.
@@
Some comments about the rules (18) and (19) are in order. The relation (18a) indicates that the inertia parameter related to the subgroup $\mathcal{P}_{1}$ increases (or decreases) when the difference between the local densities of the subgroups $\mathcal{P}_{3}$ and $\mathcal{P}_{2}$ in the subinterval of length $\tau$ increases (or decreases). This rule suggests that an increase in the cell $\alpha$ of this difference induces ordinary citizens to lower their tendency to change (police local density is greater than drug addicts one, so ordinary citizens would feel safe), whereas a decrease causes ordinary citizens to be less conservative (drug addicts local density is greater than police one, so ordinary citizens would not feel safe).
@@
The relation (18b) is similar to rule (18a): the inertia parameter related to the subgroup $\mathcal{P}_{2}$ increases (or decreases) as the difference between local densities of the subgroups $\mathcal{P}_{1}$ and $\mathcal{P}_{3}$ increases (or decreases). This rule takes into account that, when the above difference increases, the drug addicts are induced to remain in the same area (ordinary citizens local density is greater than police one, therefore drug addicts continue to commit crime), whereas a decrease causes drug addicts to be less conservative (police local density is greater than the ordinary citizens one, so that the drug addicts will tend to move to other areas to commit crimes).
@@
The rule for inertia parameters associated to the third subgroup $\mathcal{P}_{3}$, as expressed in (18c), incorporates the concept of social alarm introduced by Epstein [1]. The parameters update is influenced by the sum of interactions variation between ordinary citizens and drug addicts, alongside police local density. This social alarm is crucial for illustrating scenarios where law enforcement intervention is necessary to restore order.
@@
The relation (19a) indicates that the diffusion parameter related to ordinary citizens $\mathcal{P}_{1}$ increases (or decreases) when the difference between drug addicts and police local densities increases (or decreases). The idea behind it is that ordinary citizens must diffuse to neighboring territories when there is a high crime rate (the difference is positive), but if the law enforcement personnel is sufficient in number to maintain order (the difference is negative), ordinary citizens can remain in their safe place.
@@
The relation (19b) shows that the diffusion parameter associated to drug addicts $\mathcal{P}_{2}$ increases (or decreases) as the difference between police and ordinary citizens local densities increases (or decreases). Drug addicts are less likely to dissolve when ordinary citizens density is high and police one is low, as this encourages continued criminal activity. In fact, with a high police presence, drug addicts may relocate to seek opportunities to commit crimes elsewhere.
@@
Finally, the relation (19c) reveals that the diffusion parameter related to police $\mathcal{P}_{3}$ increases (or decreases) when the interactions between the drug addicts and ordinary citizens decrease (or increase). This implies that police should spread to other cells when social alarm in the cell $\alpha$ is low.
@@ 4 Numerical results
In this Section, we exhibit and discuss some numerical simulations for the operatorial model (11). In the first part, we neglect spatial inhomogeneities, i.e., all the interactions occur just in a single cell, whereas, the second one concerns with the model investigated in a one–dimensional spatial domain made of $N=50$ cells. In both parts, the numerical results obtained following the standard Heisenberg view and by means of the ($\mathcal{H},\rho$)–induced dynamics are compared.
@@ 4.1 Model in one cell
In this case, the time evolution is investigated in the situation where all the population is confined in only one cell: as the diffusion effects and the nonlocal competitive interactions are absent, the Hamiltonian operator reduces to the form $\mathcal{H}=\mathcal{H}_{0}+\mathcal{H}_{I}$.
In the standard Heisenberg dynamics, we see in Figures 2, 3 and 4 that the evolution exhibits a never ending oscillatory trend, as expected. On the contrary, when the rule (18) is used, the amplitude of the oscillations is decreasing in time, and the mean densities of the three subgroups tend to approach an asymptotic equilibrium (see Figures 2, 3 and 4). Furthermore, from Figures 2, 3 and 4, an interpretation can be the following: law enforcement personnel is able to better keep drug users/dealers under control. We remark that the time required to reach a stationary trend after the transient evolution depends on the choosen parameters.
Figure 2: Time evolution, in one cell, of the mean values of the number operators related to the three subgroups. No rule (Figure 2), and rule (18) with $\tau=4$ (Figure 2). The initial conditions in both cases are: $n_{1}^{0}=0.8$, $n_{2}^{0}=0.08$, $n_{3}^{0}=0.04$. The inertia and interaction parameters are $\omega_{1}=0.5$, $\omega_{2}=0.3$, $\omega_{3}=0.2$, and $\lambda_{1}=0.04$, $\lambda_{2}=0.025$, respectively.
Figure 3: Time evolution, in one cell, of the mean values of the number operators related to the three subgroups. No rule (Figure 3), and rule (18) with $\tau=4$ (Figure 3). The initial conditions in both cases are: $n_{1}^{0}=0.8$, $n_{2}^{0}=0.08$, $n_{3}^{0}=0.04$. The inertia and interaction parameters are $\omega_{1}=0.3$, $\omega_{2}=0.1$, $\omega_{3}=0.1$, and $\lambda_{1}=0.04$, $\lambda_{2}=0.025$, respectively.
Figure 4: Time evolution, in one cell, of the mean values of the number operators related to the three subgroups. No rule (Figure 4), and rule (18) with $\tau=4$ (Figure 4). The initial conditions in both cases are: $n_{1}^{0}=0.8$, $n_{2}^{0}=0.08$, $n_{3}^{0}=0.04$. The inertia and interaction parameters are $\omega_{1}=0.2$, $\omega_{2}=0.1$, $\omega_{3}=0.3$, and $\lambda_{1}=0.04$, $\lambda_{2}=0.025$, respectively.